QUESTION FOUR

4) Suppose that you have the following demand curve. 

Q = 400 – 6P + 0.0051

Where

Q = quantity demanded,

P = price, and

I = average income


You know that the current market price is ZMW20 and average income is ZMW20,000

a)                Calculate current demand.


b)                Calculate the price elasticity of demand. 


c)                Calculate the income elasticity of demand.


1
Expert's answer
2021-09-07T19:15:15-0400

Here, it is given that the demand is the function of price and income with the price pf ZMW20 and average income of ZMW20,000.



a) Using the given information, current demand is:

"Q=400-6(20)+0.005(20,000)"

"Q=380"


b) Using the given information, price elasticity of demand would be: 

"Q=400-6P+0.005I"

"dQ\/dp * P\/Q=-6 * 20\/380"

=-0.316


Therefore, price elasticity of demand is 6, which implies demand is highly price elastic.


c) Using the given information, income elasticity of demand would be:


"Q=400-6P+0.005I"


"dQ\/dI* I\/Q=0.005 *20,000\/380"

=0.2632

Therefore, income elasticity of demand is 0.005, which implies demand is highly income inelastic.